Ecoinformatics site parent site of Partnership for Biodiversity Informatics site parent site of REAP - Home




Site Logs

2008-07-21 Mon
site: REAP01
visitor: Emily Orling

Conducted experiments to 1) ensure we know the physical location of each light sensor and 2) try to determine why qsosun is tracking but has been reporting lower than ground level values since 14:30 2008-05-30.

11:10-11:30PT - covered (blocked sun) from linear light bar inside treated area (lower on the hill than station)
11:10 - 11:50 - removed soil moisture probe inside treated area. on reinsert probe leg bent a bit, but is relatively straight now.
13:50-14:00PT - covered (blocked sun) from qsosun

14:00PT cleaned qsosun w/ alcohol and qtip

Site in ok shape. Rain gage funnel and filter-mesh found on ground nearby, difficulty clipping it back on securely so was secured with duct tape in two places. Some bird waste on solar panel cleaned off. qsosun looked clean, but was cleaned anyways as noted above. qsosun cable looked fine. Putty sealant on enclosure bottom around incoming cables is starting to flake off, future visitor should bring more. Enclosure desiccant should probably be replaced too. Screw in lightening rod clamp is rusting.

Go to top   Edit this page   More info...   Attach file...
This page last changed on 05-Sep-2008 18:07:54 PDT by uid=barseghian,o=NCEAS.